# Pasadena, CA 91125, USA

### v1

@@ 4.4.3 General Form
With this at hand, we compute the $U(1)_{A}$-refined half index given in a fermionic sum formula we propose
@@ 4.5.1 $E_{6}$
The IR trace formula for the Schur index of $(A_{1},E_{6})$ theory gives
@@
As before, from this we find 3d $U(1)^{6}$ Chern-Simons theory with level $K=b^{E_{6}}$ and the mixed Chern-Simons level between the gauge $U(1)$’s and $R$ symmetry $\rho=\{2,2,\cdots,2\}$. The theory is coupled to $12$ chiral multiplets $\Phi_{i}$ and $\widetilde{\Phi}_{i}$ for $i=1,2,\cdots,6$, all with $R$-charge 0, and their gauge charges are 1-1 under $U(1)_{i}$.
From our conjectural form of the Macdonald index in section 2.3, we identify the $U(1)_{A}$ symmetry as
@@
Indeed, upon computing the $T$-refined half index with A-twist, we recover
@@
which can be understood as be the half index of a 3d $U(1)^{4}$ Chern-Simons theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$’s and R-symmetry are $\rho=\{2,4,6,4\}$. There are 4 chiral multiplets $\Phi_{i}$ for $i=1,2,3,4$, all with trial $R$ charge 0 and gauge charge 1-1 under $U(1)_{i}$.
@@
there is $U(1)^{2}$ unbroken global symmetry $U(1)_{T_{1}}+2U(1)_{T_{2}}-U(1)_{T_{4}}$ and $U(1)_{T_{3}}+U(1)_{T_{4}}$. For any other choice of three monopole deformations, only one $U(1)$ global symmetry remains unbroken, and the surviving global symmetry always takes the form of
@@
Performing $F$-maximization with this $U(1)_{A}$ (see Fig 11), and calculating the superconformal index,
Figure 11: The absolute value of three sphere partition function for the 3d theory from (158) with changing mixing parameter $\nu$ with $U(1)_{A}=U(1)_{T_{1}}+2U(1)_{T_{2}}+3U(1)_{T_{3}}+2U(1)_{T_{4}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$
@@ 4.5.2 $E_{7}$
and the mixed Chern-Simons levels between the gauge $U(1)$’s and $R$-symmetry are $\rho=\{2,2,2,2,2,2,1,1\}$. The theory is coupled to $14$ chirals multiplets, all with trial $R$ charge 0. Their gauge charges are
@@
$\quad+V_{4}\Phi_{3}\Phi_{5}+V_{4}\widetilde{\Phi}_{3}\widetilde{\Phi}_{5}+V_{5}\Phi_{4}\Phi_{6}+V_{5}XY+V_{6}\widetilde{\Phi}_{5}+V_{7,8}\Phi_{3}.$ (166)
@@
As far as the authors are aware, there is no known explicit expression for the Macdonald index of the $(A_{1},E_{7})$ theory. Therefore provides a prediction for this theory.
@@ 4.5.3 $E_{8}$
As before, from this we find 3d $U(1)^{8}$ Chern-Simons theory with level $K=b^{E_{8}}$ and the mixed Chern-Simons level between the gauge $U(1)$’s and $R$ symmetry $\rho=\{2,2,\cdots,2\}$. There are $16$ chiral multiplets $\Phi_{i}$ and $\widetilde{\Phi}_{i}$ for $i=1,2,\cdots,8$, all with $R$ charge 0, and their gauge charges are 1-1 under $U(1)_{i}$
From our conjectural form of the Macdonald index in section 2.3, we identify the $U(1)_{A}$ symmetry as
@@
Computing the $T$-refined half index with A-twist, we recover the conjecture form of the Macdonald index
@@
The last expression can be interpreted as an A-twisted half index of the 3d $U(1)^{3}$ Chern-Simons theory with level
@@
The mixed Chern-Simons level between the gauge $U(1)$s and $R$ symmetry are $\rho=\{0,0,-2\}$. There are four chiral multiplets:
  • $\Phi_{1}$ has gauge charge 1-1 under $U(1)_{1}$ and 1-1 under $U(1)_{2}$. Its $R$ charge is $2$.
  • $\Phi_{2}$ has gauge charge 1-1 under $U(1)_{1}$ and +1+1 under $U(1)_{3}$. Its $R$ charge is 0.
  • $\Phi_{3}$ has gauge charge 1-1 under $U(1)_{2}$ and +1+1 under $U(1)_{3}$. Its $R$ charge is 0.
  • $\Phi_{4}$ has gauge charge 1-1 under $U(1)_{3}$. Its $R$ charge is $2$.
@@
Figure 12: The absolute value of three sphere partition function for the 3d theory from (175) with changing mixing parameter $\nu$ with $U(1)_{A}=U(1)_{\Phi_{2}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$
Performing the $F$-maximization with this $U(1)_{A}$ (see Fig. 12), and calculating the superconformal index, which reads,
@@
This provides a strong consistency check of our proposal.
@@ Acknowledgements.
This work is supported by the National Research Foundation of Korea (NRF) Grant RS-2024-00405629, as well as the KAIST-KIAS collaboration program. The work of Heeyeon Kim and Hongseok Kim is also supported by the National Research Foundation of Korea (NRF) Grant NRF2023R1A2C1004965 and by the POSCO Science Fellowship of POSCO TJ Park Foundation. The work of J.S. is also supported by the National Research Foundation of Korea (NRF) Grant RS-2023-00208602, and also by the Walter Burke Institute for Theoretical Physics at Caltech.

### v2

@@ 4.4.3 General Form
With this at hand, we compute the $U(1)_{A}$-refined half-index given in a fermionic sum formula we propose
@@ 4.5.1 $E_{6}$
The IR trace formula for the Schur index of the $(A_{1},E_{6})$ theory gives
@@
As before, from this we find the 3d $U(1)^{6}$ Chern-Simons theory with level $K=b^{E_{6}}$ and the mixed Chern-Simons levels between the gauge $U(1)$s and $R$-symmetry $\rho=\{2,2,\cdots,2\}$. The theory is coupled to $12$ chiral multiplets $\Phi_{i}$ and $\widetilde{\Phi}_{i}$ for $i=1,2,\cdots,6$, all with $R$-charge 0, and their gauge charges are 1-1 under $U(1)_{i}$.
From our conjectural form of the Macdonald index in section 2.3, we conjecture that the $U(1)_{A}$ symmetry can be identified with
@@
Indeed, upon computing the $T$-refined half-index with A-twist, we recover
@@
which can be understood as the half-index of a 3d $U(1)^{4}$ Chern-Simons theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$s and $R$-symmetry are $\rho=\{2,4,6,4\}$. There are 4 chiral multiplets $\Phi_{i}$ for $i=1,2,3,4$, all with trial $R$-charge 0 and gauge charge 1-1 under $U(1)_{i}$.
@@
there remains $U(1)^{2}$ unbroken global symmetry $U(1)_{T_{1}}+2U(1)_{T_{2}}-U(1)_{T_{4}}$ and $U(1)_{T_{3}}+U(1)_{T_{4}}$. For any other choice of three monopole deformations, only one $U(1)$ global symmetry remains unbroken, and the surviving global symmetry always takes the form of
@@
Performing $F$-maximization with this $U(1)_{A}$ (see Figure 11), and calculating the superconformal index,
Figure 11: The absolute value of three-sphere partition function for the 3d theory from (158) as a function of the mixing parameter $\nu$ with $U(1)_{A}=U(1)_{T_{1}}+2U(1)_{T_{2}}+3U(1)_{T_{3}}+2U(1)_{T_{4}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$
@@ 4.5.2 $E_{7}$
and the mixed Chern-Simons levels between the gauge $U(1)$s and the $R$-symmetry are $\rho=\{2,2,2,2,2,2,1,1\}$. The theory is coupled to $14$ chiral multiplets, all with trial $R$-charge 0. Their gauge charges are
@@
$\qquad+V_{4}\Phi_{3}\Phi_{5}+V_{4}\widetilde{\Phi}_{3}\widetilde{\Phi}_{5}+V_{5}\Phi_{4}\Phi_{6}+V_{5}XY+V_{6}\widetilde{\Phi}_{5}+V_{7,8}\Phi_{3}.$ (166)
@@
As far as the authors are aware, there is no known explicit expression for the Macdonald index of the $(A_{1},E_{7})$ theory. This therefore provides a prediction for this theory.
@@ 4.5.3 $E_{8}$
As before, from this we find the 3d $U(1)^{8}$ Chern-Simons theory with level $K=b^{E_{8}}$ and the mixed Chern-Simons levels between the gauge $U(1)$s and $R$-symmetry $\rho=\{2,2,\cdots,2\}$. There are $16$ chiral multiplets $\Phi_{i}$ and $\widetilde{\Phi}_{i}$ for $i=1,2,\cdots,8$, all with $R$-charge 0, and their gauge charges are 1-1 under $U(1)_{i}$
From our conjectural form of the Macdonald index in section 2.3, we conjecture that the $U(1)_{A}$ symmetry is identified with
@@
Computing the $T$-refined half-index with A-twist, we recover the conjecture form of the Macdonald index
@@
The last expression can be interpreted as an A-twisted half-index of the 3d $U(1)^{3}$ Chern-Simons theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$s and the $R$-symmetry are $\rho=\{0,0,-2\}$. There are four chiral multiplets:
  • $\Phi_{1}$ has gauge charge 1-1 under $U(1)_{1}$ and 1-1 under $U(1)_{2}$. Its $R$-charge is $2$.
  • $\Phi_{2}$ has gauge charge 1-1 under $U(1)_{1}$ and +1+1 under $U(1)_{3}$. Its $R$-charge is 0.
  • $\Phi_{3}$ has gauge charge 1-1 under $U(1)_{2}$ and +1+1 under $U(1)_{3}$. Its $R$-charge is 0.
  • $\Phi_{4}$ has gauge charge 1-1 under $U(1)_{3}$. Its $R$-charge is $2$.
@@
Figure 12: The absolute value of three-sphere partition function for the 3d theory from (175) as a function of the mixing parameter $\nu$ with $U(1)_{A}=U(1)_{\Phi_{2}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$
The $F$-maximization with this $U(1)_{A}$ (see Figure 12), and the superconformal index, which reads,
@@
provide a strong consistency check of our proposal.
@@ Acknowledgements.
This work is supported by the National Research Foundation of Korea (NRF) Grant RS-2024-00405629, as well as the KAIST-KIAS collaboration program. The work of Heeyeon Kim and Hongseok Kim is also supported by the National Research Foundation of Korea (NRF) Grant NRF2023R1A2C1004965 and by the POSCO Science Fellowship of POSCO TJ Park Foundation. The work of J.S. is also supported by the National Research Foundation of Korea (NRF) Grant RS-2023-00208602, and also in part by the Walter Burke Institute for Theoretical Physics at Caltech and by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632.